OpenStudy (anonymous):

[8.01] Factor completely: −5x2 + 10x − 15 Can someone show me how to do this please?

OpenStudy (anonymous):

Factor your back numbers first

OpenStudy (anonymous):

*number

OpenStudy (anonymous):

back number?

OpenStudy (anonymous):

15 would be your back number.

OpenStudy (anonymous):

ok , after I do that, whats next?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

help with this

OpenStudy (anonymous):

\[-5x ^{2}+10x-15=-5\left( x ^{2} -2x\right)-15=-5\left( x ^{2} -2x+1-1\right)-15\] \[=-5\left( x-1 \right)^{2}+5-15=-5\left\{ \left( x-1 \right) ^{2}+2\right\}\] \[=-5\left\{ \left( x-1 \right)^{2} -\left( -2 \right)\right\}\] \[=-5\left\{ \left( x-1 \right)^{2} -\left( \iota \sqrt{2} \right)^{2}\right\}\] \[=-5\left( x-1+\iota \sqrt{2} \right)\left( x-1-\iota \sqrt{2} \right)\]